Could you please help me write an equation in standard form that satisfies these following conditions:
Parallel to y=4x+1 and passes through (-3, 5).
Please show me how, step-by-step!!
Thank you :)

Found 2 solutions by sudhanshu_kmr, htmentor:
Answer by sudhanshu_kmr(1152)


Parallel to y=4x+1 and passes through (-3, 5).

equation of parallel line is : y= 4x + c
where c is any real value.
to find the value of c, put the value of x and y.
5 = 4(-3) + c
=> c = 17

equation of line is : y = 4x + 17

Answer by htmentor(1343)

Parallel lines have the same slope. Therefore, the slope = 4. Now we need to find the y-intercept.
The slope of a line is determined from two points on the line (x1,y1), (x2,y2):
m = (y2-y1)/(x2-x1) [rise/run]
One of the points on the line is the y-intercept with coordinates (0,b), and the other point is (-3,5).
Use the formula for the slope to find b:
m = 4 = (b-5)/(0+3) -> b-5 = 12, or b = 17
So the equation for the line is:
y = 4x + 17